# Online Detecting LLM-Generated Texts via Sequential Hypothesis Testing by Betting

### v1

@@ 4.2 Experimental Results
Comparisons with Baselines. In this part, we use the score function of Fast-DetectGPT and scoring model Neo-2.7 to get text scores, and then compare the performance of our method with two baselines that adapt the fixed-time permutation test to the sequential hypothesis setting. Batch sizes $k\in\{25,50,100,250,500\}$ are considered for the baselines. We set the estimated $\epsilon$ and $d_{*}$ values the same as in Scenario 2. The baselines are also implemented in a manner to conduct the composite hypothesis test. We observe a significant difference between the scores of XSum texts and machine-generated news, which causes the baselines of the permutation test to reject the null hypothesis most of the time immediately after receiving the first batch. This in turn results in the nearly vertical lines observed in the left subfigure of Figure 4(a) and Figure 4(b), where the averaged rejection time across 1000 repeated tests closely approximates the batch size $k$ for each of the 20 significance levels $\alpha$. On the other hand, we observe that when $H_{0}$ is true, the baselines might not be a valid $\alpha$-test, even with a corrected significance level. This arises from the increased variability of text scores introduced by smaller batch sizes, which results in observed absolute differences in means that may exceed the $\epsilon$ value under $H_{0}$.
Our method can quickly detect an LLM while keeping the false positive rates (FPRs) below the specified significance levels, which is a delicate balance that can be difficult for the baselines to achieve. Without prior knowledge of the $\epsilon$ value, the baselines of permutation tests may fail to control the type-I error with small batch sizes and cannot quickly reject the null hypothesis while ensuring that FPRs remain below $\alpha$, unlike our method.
5 Limitations and Outlooks
@@
In this paper, we demonstrate that our algorithm, which builds on the score functions of offline detectors, can rapidly determine whether a stream of text is generated by an LLM and provides strong statistical guarantees. Specifically, it controls the type-I error rate below the significance level, ensures that the source of LLM-generated texts can eventually be identified, and guarantees an upper bound on the expected time to correctly declare the unknown source as an LLM under a mild assumption. Although the choice of detector can influence the algorithm’s performance and some parameters related to text scores need to be predefined before receiving the text, our experimental results show that most existing detectors provide effective score functions, and our method performs well in most cases when using estimated values of parameters based on text scores from the first few time steps. To further enhance its efficacy, it may be worthwhile to design score functions tailored to the sequential setting, improve parameter estimations with scores from more time steps, and study the trade-offs between delaying the start of testing and obtaining more reliable estimates.
Moreover, our algorithm could potentially be used as an effective tool to promptly identify and mitigate the misuse of LLMs in generating texts, such as identifying users who post LLM-generated comments in forums to manipulate public opinion, monitoring social media accounts that disseminate harmful information generated by LLMs, and rapidly detecting sources of fake news generated by LLMs on public websites. Exploring these applications could be a promising direction.
@@ A.1 Related works of detecting machine-generated texts
Some methods distinguish between human-written and machine-generated texts by comparing their statistical properties[Jawahar et al., 2020]. Lavergne et al. [2008] introduced a method, which uses relative entropy scoring to effectively identify texts from Markovian text generators. Perplexity is also a metric for detection, which quantifies the uncertainty of a model in predicting text sequences[Hashimoto et al., 2019]. Gehrmann et al. [2019] developed GLTR tool, which leverages statistical features such as per-word probability, rank, and entropy, to enhance the accuracy of fake-text detection. Mitchell et al. [2023] created a novel detector called DetectGPT, which identifies a machine-generated text by noting that it will exhibit higher log-probability than samples where some words of the original text have been rewritten/perturbed. Su et al. [2023] then introduced two advanced methods utilizing two metrics: Log-Likelihood Log-Rank Ratio (LRR) and Normalized Perturbed Log Rank (NPR), respectively. Bao et al. [2023] developed Fast-DetectGPT, which replaces the perturbation step of DetectGPT with a more efficient sampling operation. Solaiman et al. [2019] employed the classic logistic regression model on TF-IDF vectors to detect texts generated by GPT-2, and noted that texts from larger GPT-2 models are more challenging to detect than those from smaller GPT-2 models. Researchers have also trained supervised models on neural network bases. Bakhtin et al. [2019] found that Energy-based models (EBMs) outperform the behavior of using the original language model log-likelihood in real and fake text discrimination. Zellers et al. [2019] developed a robust detection method named GROVER by using a linear classifier, which can effectively spot AI-generated ‘neural’ fake news. Ippolito et al. [2019] showed that BERT-based[Devlin, 2018] classifiers outperform humans in identifying texts characterized by statistical anomalies, such as those where only the top k high-likelihood words are generated, yet humans excel at semantic understanding. Solaiman et al. [2019] fine-tuned RoBERTa[Liu et al., 2019] on GPT-2 outputs and achieved approximately $95\%$ accuracy in detecting texts generated by 1.5 billion parameter GPT-2. The effectiveness of RoBERTa-based detectors is further validated across different text types, including machine-generated tweets[Fagni et al., 2021], news articles[Uchendu et al., 2020], and product reviews[Adelani et al., 2020]. Other supervised classifiers, such as GPTZero[Tian, 2023] and OpenAI’s Classifier[OpenAI, 2023], have also proven to be strong detectors. Moreover, some research has explored watermarking methods that embed detectable patterns in LLM-generated texts for identifying, see e.g., Jalil and Mirza [2009], Kamaruddin et al. [2018], Abdelnabi and Fritz [2021], Zhao et al. [2023], Kirchenbauer et al. [2023], Christ et al. [2024]
@@ A.2 Related Works of sequential hypothesis testing by betting
Kelly [1956] first proposed a strategy for sequential betting with initial wealth on the outcome of each coin flip $g_{t}$ in round $t$, which can take values of $+1$ (head) or $-1$ (tail), generated i.i.d. with the probability of heads $p\in[0,1]$. It is shown that betting a fraction $\beta_{t}=2p-1$ on heads in each round will yield more accumulated wealth than betting any other fixed fraction of the current wealth in the long run. Orabona and Pál [2016] demonstrated the equivalence between the minimum wealth of betting and the maximum regret in one-dimensional Online Linear Optimization (OLO) algorithms, which introduces the coin-betting abstraction for the design of parameter-free algorithms. Based on this foundation, Cutkosky and Orabona [2018] developed a coin betting algorithm, which uses an exp-concave optimization approach through the Online Newton Step (ONS). Subsequently, Shekhar and Ramdas [2023] applied their betting strategy, along with the general principles of testing by betting as clarified by Shafer [2021], to nonparametric two-sample hypothesis testing. Chugg et al. [2023] then conducted sequentially audits on both classifiers and regressors within the general two-sample testing framework established by Shekhar and Ramdas [2023], which demonstrate that this method remains robust even in the face of distribution shifts. Additionally, other studies[Orabona and Jun, 2023, Waudby-Smith and Ramdas, 2024] have developed practical strategies that leverage online convex optimization methods, with which the betting fraction can be adaptively selected to provide statistical guarantees for the results.
@@ Appendix B Related Score Functions
where $\tilde{\mu}=\frac{1}{m}\sum_{i=1}^{m}\log p_{\theta}(\tilde{x}^{(i)})$, and $\tilde{\sigma}^{2}=\frac{1}{m}\sum_{i=1}^{m}\left[\log p_{\theta}(\tilde{x}^{(i)})-\tilde{\mu}\right]^{2}$. We can write $\log p_{\theta}(x)$ as $\sum_{j=1}^{n}\log p_{\theta}(x_{j}|x_{1:j-1})$, where $n$ denotes the number of tokens of $x$, $x_{j}$ denotes the $j$-th token, and $x_{1:j-1}$ means the first $(j-1)$ tokens. Similarly, $\log p_{\theta}(\tilde{x}^{(i)})=\sum_{j=1}^{\tilde{n}^{(i)}}\log p_{\theta}(\tilde{x}_{j}^{(i)}|\tilde{x}_{1:j-1}^{(i)})$, where $\tilde{n}^{(i)}$ is the number of tokens of $i$-th perturbed sample $\tilde{x}^{(i)}$.
@@
where $\tilde{x}_{j}^{(i)}$ denotes the $i$-th generated sample for the $j_{th}$ token of the text $x$, $q_{\zeta}(\tilde{x}_{j}^{(i)}|x_{1:j-1})$ is the probability of this sampled token given by the sampling model according to the probability distribution of all possible tokens at position $j$, conditioned on the first $(j-1)$ tokens of $x$. Besides, $p_{\theta}(\tilde{x}_{j}^{(i)}|x_{1:j-1})$ is the conditional probability of $\tilde{x}_{j}^{(i)}$ evaluated by the scoring model, $s$ represents the total number of possible tokens at each position, corresponding to the size of the entire vocabulary used by the sampling model. We can use the same value at each position in the formula, as the vocabulary size remains consistent across positions. The sample mean and the variance can be considered in practice
@@ Appendix C Proof of Regret Bound of ONS
Under $H_{0}$, i.e., $\mu_{x}=\mu_{y}$, the wealth is a P-supermartingale. The value of $g_{t}$ are constrained to the interval $[-1,1]$ in previous works[Orabona and Pál, 2016, Cutkosky and Orabona, 2018, Chugg et al., 2023] with the scores $\phi({x}_{t})$ and $\phi({y}_{t})$ each range from $[0,1]$. To ensure that wealth $W_{t}$ remains nonnegative and to establish the regret bound by ONS, $\theta_{t}$ is always selected within $[-1/2,1/2]$. In our setting, however, the actual range of score difference between two texts, denoted as $g_{t}=\phi({x}_{t})-\phi({y}_{t})$, is typically unknown beforehand. If we assume that the ranges for both $\phi({x}_{t})$ and $\phi({y}_{t})$ are $[m_{t},n_{t}]$, the range of their difference $g_{t}$ is then symmetric about zero, which spans from $-(n_{t}-m_{t})$ to $(n_{t}-m_{t})$. we suppose an upper bound value $d_{t}\geq n_{t}-m_{t}$ and express the range as $g_{t}\in[-d_{t},d_{t}]$, where $d_{t}\geq 0$. Then, choosing $\theta_{t}$ within $[-1/2d_{t},1/2d_{t}]$ can guarantee that the wealth is a nonnegative P-supermartingale. If we consider the condition of either $W_{t}\geq 1/\alpha\text{ or }W_{T}\geq Z/\alpha$ for any stopping time $T$ as the indication to "reject $H_{0}$" and apply the randomized Ville’s inequality[Ramdas and Manole, 2023], the type-I error can be controlled below the significance level $\alpha$ under $H_{0}$.
@@
where $G_{t}\cdot D_{t}$ is a constant for all $t$, and $z_{t}$, $a_{t}$, and $\mathcal{K}_{t}$ are defined in Algorithm 2.
@@
We recall that $D_{t}$ is defined as the diameter of $\mathcal{K}_{t}$, i.e., $D_{t}=\max_{a,b\in\mathcal{K}_{t}}\lVert a-b\rVert$, and $G_{t}$ is defined as a bound on the gradients of the loss function $\ell_{t}(\theta)=-\ln(1-g_{t}\theta)$ at time $t$, i.e., $G_{t}=\max_{\theta_{t}\in\mathcal{K}_{t}}\lvert\frac{d}{d\theta_{t}}\ell_{t}(\theta_{t})\rvert$. In our setting, $\mathcal{K}_{t}$ is $[-{1}/{2d_{t}},{1}/{2d_{t}}]$, thus $D_{t}={1}/{d_{t}}$. The gradient $z_{t}=\nabla\ell_{t}(\theta_{t})={g_{t}}/{(1-g_{t}\theta_{t})}$. We find that $\ell_{t}$ monotonically increases with $g_{t}$ and $\theta_{t}$, which means $G_{t}$ can be taken at the maximum $g_{t}$ and the maximum $\theta_{t}$. Since $g_{t}=\phi({x}_{t})-\phi({y}_{t})\in[-d_{t},d_{t}]$, we have $G_{t}={d_{t}}/{(1-d_{t}\cdot\frac{1}{2d_{t}})}=2d_{t}$. Above all, we get $G_{t}\cdot D_{t}=2d_{t}\cdot 1/d_{t}=2$ for each $t$, and $\alpha=1$. The value of $\gamma_{t}=\frac{1}{2}\min\{{1}/{G_{t}D_{t}},\alpha\}$ becomes a fixed positive constant for all $t$. Therefore, we can simply use $\gamma$ in the remaining proof since $\gamma_{t}$ is the same for every $t$.

### v2

@@ 4 Experiments
Comparisons with Baselines. In this part, we use the score function of Fast-DetectGPT and scoring model Neo-2.7 to get text scores, and then compare the performance of our method with two baselines that adapt the fixed-time permutation test to the sequential hypothesis setting. Batch sizes $k\in\{25,50,100,250,500\}$ are considered for the baselines. We set the estimated $\epsilon$ and $d_{*}$ values the same as in Scenario 2. The baselines are also implemented in a manner to conduct the composite hypothesis test. We observe a significant difference between the scores of XSum texts and machine-generated news, which causes the baselines of the permutation test to reject the null hypothesis most of the time immediately after receiving the first batch. This in turn results in the nearly vertical lines observed in the left subfigure of Figure 4(a) and Figure 4(b), where the averaged rejection time across 1000 repeated tests closely approximates the batch size $k$ for each of the 20 significance levels $\alpha$. On the other hand, we observe that when $H_{0}$ is true, the baselines might not be a valid $\alpha$-test, even with a corrected significance level. This arises from the increased variability of text scores introduced by smaller batch sizes, which results in observed absolute differences in means that may exceed the $\epsilon$ value under $H_{0}$. Our method can quickly detect an LLM while keeping the false positive rates (FPRs) below the specified significance levels, which is a delicate balance that can be difficult for the baselines to achieve. Without prior knowledge of the $\epsilon$ value, the baselines of permutation tests may fail to control the type-I error with small batch sizes and cannot quickly reject the null hypothesis while ensuring that FPRs remain below $\alpha$, unlike our method.
Extensions. We further consider (i) an experiment where the sequential texts are produced by various LLMs instead of a single LLM, and (ii) a scenario that the unknown source publishes both human-written and LLM-generated texts, where the null hypothesis becomes that all texts from the unknown source are human-written, and the alternative hypothesis posits that not all texts are human-written. We refer the reader to Appendix I for the encouraging results.
5 Conclusion
@@
In this paper, we demonstrate that our algorithm, which builds on the score functions of offline detectors, can rapidly determine whether a stream of text is generated by an LLM and provides strong statistical guarantees. Specifically, it controls the type-I error rate below the significance level, ensures that the source of LLM-generated texts can eventually be identified, and guarantees an upper bound on the expected time to correctly declare the unknown source as an LLM under a mild assumption. Although the choice of detector can influence the algorithm’s performance and some parameters related to text scores need to be predefined before receiving the text, our experimental results show that most existing detectors provide effective score functions, and our method performs well in most cases when using estimated values of parameters based on text scores from the first few time steps. To further enhance its efficacy, it may be worthwhile to design score functions tailored to the sequential setting, improve parameter estimations with scores from more time steps, and study the trade-offs between delaying the start of testing and obtaining more reliable estimates. Moreover, our algorithm could potentially be used as an effective tool to promptly identify and mitigate the misuse of LLMs in generating texts, such as monitoring social media accounts that disseminate harmful information generated by LLMs, rapidly detecting sources of fake news generated by LLMs on public websites, and identifying users who post LLM-generated comments to manipulate public opinion. Exploring these applications could be a promising direction.
Acknowledgments
The authors gratefully acknowledge the support provided by the Gemma Academic Program and Google Cloud Credits.
@@ A.1 Related works of detecting machine-generated texts
Some methods distinguish between human-written and machine-generated texts by comparing their statistical properties [Jawahar et al., 2020]. Lavergne et al. [2008] introduced a method, which uses relative entropy scoring to effectively identify texts from Markovian text generators. Perplexity is also a metric for detection, which quantifies the uncertainty of a model in predicting text sequences [Hashimoto et al., 2019]. Gehrmann et al. [2019] developed GLTR tool, which leverages statistical features such as per-word probability, rank, and entropy, to enhance the accuracy of fake-text detection. Mitchell et al. [2023] created a novel detector called DetectGPT, which identifies a machine-generated text by noting that it will exhibit higher log-probability than samples where some words of the original text have been rewritten/perturbed. Su et al. [2023] then introduced two advanced methods utilizing two metrics: Log-Likelihood Log-Rank Ratio (LRR) and Normalized Perturbed Log Rank (NPR), respectively. Bao et al. [2023] developed Fast-DetectGPT, which replaces the perturbation step of DetectGPT with a more efficient sampling operation. Solaiman et al. [2019] employed the classic logistic regression model on TF-IDF vectors to detect texts generated by GPT-2, and noted that texts from larger GPT-2 models are more challenging to detect than those from smaller GPT-2 models. Researchers have also trained supervised models on neural network bases. Bakhtin et al. [2019] found that Energy-based models (EBMs) outperform the behavior of using the original language model log-likelihood in real and fake text discrimination. Zellers et al. [2019] developed a robust detection method named GROVER by using a linear classifier, which can effectively spot AI-generated ‘neural’ fake news. Ippolito et al. [2019] showed that BERT-based [Devlin, 2018] classifiers outperform humans in identifying texts characterized by statistical anomalies, such as those where only the top k high-likelihood words are generated, yet humans excel at semantic understanding. Solaiman et al. [2019] fine-tuned RoBERTa [Liu et al., 2019] on GPT-2 outputs and achieved approximately $95\%$ accuracy in detecting texts generated by 1.5 billion parameter GPT-2. The effectiveness of RoBERTa-based detectors is further validated across different text types, including machine-generated tweets [Fagni et al., 2021], news articles [Uchendu et al., 2020], and product reviews [Adelani et al., 2020]. Other supervised classifiers, such as GPTZero [Tian, 2023] and OpenAI’s Classifier [OpenAI, 2023], have also proven to be strong detectors. Moreover, some research has explored watermarking methods that embed detectable patterns in LLM-generated texts for identifying, see e.g., Jalil and Mirza [2009], Kamaruddin et al. [2018], Abdelnabi and Fritz [2021], Zhao et al. [2023], Kirchenbauer et al. [2023], Christ et al. [2024]. Recently, Kobak et al. [2024] introduced “excess word usage”, a novel data-driven approach that identifies LLM usage in academic writing and avoids biases that could be potentially introduced by generation prompts from traditional human text datasets.
@@ A.2 Related Works of sequential hypothesis testing by betting
Kelly [1956] first proposed a strategy for sequential betting with initial wealth on the outcome of each coin flip $g_{t}$ in round $t$, which can take values of $+1$ (head) or $-1$ (tail), generated i.i.d. with the probability of heads $p\in[0,1]$. It is shown that betting a fraction $\beta_{t}=2p-1$ on heads in each round will yield more accumulated wealth than betting any other fixed fraction of the current wealth in the long run. Orabona and Pál [2016] demonstrated the equivalence between the minimum wealth of betting and the maximum regret in one-dimensional Online Linear Optimization (OLO) algorithms, which introduces the coin-betting abstraction for the design of parameter-free algorithms. Based on this foundation, Cutkosky and Orabona [2018] developed a coin betting algorithm, which uses an exp-concave optimization approach through the Online Newton Step (ONS). Subsequently, Shekhar and Ramdas [2023] applied their betting strategy, along with the general principles of testing by betting as clarified by Shafer [2021], to nonparametric two-sample hypothesis testing. Chugg et al. [2023] then conducted sequentially audits on both classifiers and regressors within the general two-sample testing framework established by Shekhar and Ramdas [2023], which demonstrate that this method remains robust even in the face of distribution shifts. Additionally, other studies [Orabona and Jun, 2023, Waudby-Smith and Ramdas, 2024] have developed practical strategies that leverage online convex optimization methods, with which the betting fraction can be adaptively selected to provide statistical guarantees for the results. We also note that sequential hypothesis testing by betting has drawn significant attention in recent years, e.g., Vovk and Wang [2021], Shaer et al. [2023], Ramdas et al. [2023], Podkopaev and Ramdas [2023], Pandeva et al. [2024], Podkopaev et al. [2023], Teneggi and Sulam [2024], Grünwald et al. [2024], Podkopaev et al. [2024], Cho et al. [2024], Fischer and Ramdas [2024], Bar et al. [2024].
@@ Appendix B Related Score Functions
where $\tilde{\mu}=\frac{1}{m}\sum_{i=1}^{m}\log p_{\theta}(\tilde{x}^{(i)})$, and $\tilde{\sigma}^{2}=\frac{1}{m}\sum_{i=1}^{m}\left[\log p_{\theta}(\tilde{x}^{(i)})-\tilde{\mu}\right]^{2}$. We can write $\log p_{\theta}(x)$ as $\sum_{j=1}^{n}\log p_{\theta}(x_{j}|x_{1:j-1})$, where $n$ denotes the number of tokens of $x$, $x_{j}$ denotes the $j$-th token, and $x_{1:j-1}$ means the first $j-1$ tokens. Similarly, $\log p_{\theta}(\tilde{x}^{(i)})=\sum_{j=1}^{\tilde{n}^{(i)}}\log p_{\theta}(\tilde{x}_{j}^{(i)}|\tilde{x}_{1:j-1}^{(i)})$, where $\tilde{n}^{(i)}$ is the number of tokens of $i$-th perturbed sample $\tilde{x}^{(i)}$.
@@
where $\tilde{x}_{j}^{(i)}$ denotes the $i$-th generated sample for the $j_{th}$ token of the text $x$, $q_{\zeta}(\tilde{x}_{j}^{(i)}|x_{1:j-1})$ is the probability of this sampled token given by the sampling model according to the probability distribution of all possible tokens at position $j$, conditioned on the first $j-1$ tokens of $x$. Besides, $p_{\theta}(\tilde{x}_{j}^{(i)}|x_{1:j-1})$ is the conditional probability of $\tilde{x}_{j}^{(i)}$ evaluated by the scoring model, $s$ represents the total number of possible tokens at each position, corresponding to the size of the entire vocabulary used by the sampling model. We can use the same value at each position in the formula, as the vocabulary size remains consistent across positions. The sample mean and the variance can be considered in practice
@@ Appendix C Proof of Regret Bound of ONS
Under $H_{0}$, i.e., $\mu_{x}=\mu_{y}$, the wealth is a P-supermartingale. The value of $g_{t}$ are constrained to the interval $[-1,1]$ in previous works [Orabona and Pál, 2016, Cutkosky and Orabona, 2018, Chugg et al., 2023] with the scores $\phi({x}_{t})$ and $\phi({y}_{t})$ each range from $[0,1]$. To ensure that wealth $W_{t}$ remains nonnegative and to establish the regret bound by ONS, $\theta_{t}$ is always selected within $[-1/2,1/2]$. In our setting, however, the actual range of score difference between two texts, denoted as $g_{t}=\phi({x}_{t})-\phi({y}_{t})$, is typically unknown beforehand. If we assume that the ranges for both $\phi({x}_{t})$ and $\phi({y}_{t})$ are $[m_{t},n_{t}]$, the range of their difference $g_{t}$ is then symmetric about zero, which spans from $-(n_{t}-m_{t})$ to $(n_{t}-m_{t})$. we denote an upper bound $d_{t}\geq n_{t}-m_{t}$ and express the range as $g_{t}\in[-d_{t},d_{t}]$, where $d_{t}\geq 0$. Then, choosing $\theta_{t}$ within $[-1/2d_{t},1/2d_{t}]$ can guarantee that the wealth is a nonnegative P-supermartingale. If we consider the condition of either $W_{t}\geq 1/\alpha\text{ or }W_{T}\geq Z/\alpha$ for any stopping time $T$ as the indication to "reject $H_{0}$" and apply the randomized Ville’s inequality [Ramdas and Manole, 2023], the type-I error can be controlled below the significance level $\alpha$ under $H_{0}$.
@@
where $G_{t}D_{t}$ is a constant for all $t$, and $z_{t}$, $a_{t}$, and $\mathcal{K}_{t}$ are defined in Algorithm 2.
@@
We recall that $D_{t}$ is defined as the diameter of $\mathcal{K}_{t}$, i.e., $D_{t}=\max_{a,b\in\mathcal{K}_{t}}\lVert a-b\rVert$, and $G_{t}$ is defined as a bound on the gradients of the loss function $\ell_{t}(\theta)=-\ln(1-g_{t}\theta)$ at time $t$, i.e., $G_{t}=\max_{\theta_{t}\in\mathcal{K}_{t}}\lvert\frac{d}{d\theta_{t}}\ell_{t}(\theta_{t})\rvert$. In our setting, $\mathcal{K}_{t}$ is $[-{1}/{2d_{t}},{1}/{2d_{t}}]$, thus $D_{t}={1}/{d_{t}}$. The gradient $z_{t}=\nabla\ell_{t}(\theta_{t})={g_{t}}/{(1-g_{t}\theta_{t})}$. We find that $\ell_{t}$ monotonically increases with $g_{t}$ and $\theta_{t}$, which means $G_{t}$ can be taken at the maximum $g_{t}$ and the maximum $\theta_{t}$. Since $g_{t}=\phi({x}_{t})-\phi({y}_{t})\in[-d_{t},d_{t}]$, we have $G_{t}={d_{t}}/{(1-d_{t}\cdot\frac{1}{2d_{t}})}=2d_{t}$. Above all, we get $G_{t}D_{t}=2d_{t}\cdot 1/d_{t}=2$ for each $t$, and $\alpha=1$. The value of $\gamma_{t}=\frac{1}{2}\min\{{1}/{G_{t}D_{t}},\alpha\}$ remains a positive constant for all $t$. Therefore, we can simply use $\gamma$ in the following proof since $\gamma_{t}$ is the same for all $t$.